Subject: APO 3.5-m users committee minutes, 4/13/09
From: Suzanne Hawley
Submitted: Fri, 17 Apr 2009 16:29:58 -0700 (PDT)
Message number: 1130
(previous: 1129,
next: 1131
up: Index)
APO 3.5-m Users Committee Phonecon, 4/13/09
Attending: Suzanne Hawley, Scott Anderson, Mark Klaene, Russet McMillan, Bill
Ketzeback, John Bally, Bruce Gillespie, Michael Strauss, Jon Holtzman, Remy
Indebetouw, and guests Cynthia Froning and Eric Burgh
Absent: Al Harper, Jon Fulbright
**********************************
User feedback, comments from institutional representatives:
Princeton (Michael Strauss) - Michael felt that things were going well for PU
users, but noted the recent trouble with guiding during TripleSpec use. Mark
noted that the site staff are looking into the issue, which appears to be a
known problem with the off-balance configuration of the instrument due to its
asymmetrical LN2 storage system, which is difficult to overcome with hardware.
Michael suggested that if there are predictable instrument orientations that
cause problems, perhaps these could be communicated to the users and included
in the documentation. Monitoring the fill status of the LN2 tank is difficult
to do, and it seems that there is not yet a clear discernible pattern of fill
condition vs. orientation that causes problems. Suzanne mentioned that we
could revert to operating the instrument in constrained rotation angles.
Russet added that we do not currently have the software tools, which are
non-trivial, to predict the new rotator angle before the telescope slew command
is issued.
Colorado (John Bally) - John said that he had heard no feedback from CU users.
Johns Hopkins - no report.
Washington (Scott Anderson) - Scott had nothing new to report.
Chicago - no report.
New Mexico State (Jon Holtzman) - Jon had nothing new to report.
Virginia (Remy Indebetouw) - Remy said that he had no user comments, and asked
if he needed to solicit more UVa user input on the instrument upgrade studies,
to which the answer is yes (see below).
**********************************
Discussion of telescope/instruments report:
Regarding the more detailed report below, Mark highlighted the onset of the
annual spring winds and dust. The 3.5-m telescope participated in the recent
100-hour International Year of Astronomy webcast, and the observatory hosted an
evening open house for some 90 people who were given tours of the telescope and
views of the sky through the 3.5-m telescope and ARCSAT, although the weather
was largely uncooperative. Michael asked where the open-house participants
were from, and Mark said that most were from Otero County and west Texas,
although one couple had come from Illinois to see APO and the telescopes.
The telescope and instruments have been trouble-free during the period. The
new NIC-FPS fan-out board and the NIR etalon are soon to be tested at CU. The
array power issue on TripleSpec is under study, using spare power supplies and
controllers. Agile is unchanged, and still available at NA2. APOLLO is
running well, and its new gravimeter has been installed in the cone room under
the telescope. The new direct-drive system components are being fabricated for
installation during the summer shutdown.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 3/01/09 through
4/10/09
Mark Klaene
0) Overview
Spring winds have arrived with their associated high-dust events. This has
affected a couple of nights of observing, but mostly because of the high winds
rather than elevated dust levels. We participated in the "100 hours of
Astronomy" webcast on April 3rd from the 3.5-m control room. We also held our
first-ever open house on April 10th, using a half night of engineering time for
public viewing with the 3.5- meter, and eyepiece viewing with ARCSAT and the
NMSU 1-meter telescope. 60 people had made reservations for the tours.
1) Telescope
Telescope operations have been generally smooth, with improved pointing and
tracking. No significant telescope issues arose this past month. We continue
to work on network conversion and hot- swappable backup computers.
2) Instruments
DIS, SPIcam, and Echelle are fully operational. A set of Stromgren filters
arrived and were placed in holders.
NIC-FPS has been operational with no problems. The new fan-out board is being
constructed; the new board will be tested at UVa and CU, and then a plan for
its installation will be established. We hope to integrate and test the new
board during the summer, but CU commitments to HST/COS may delay the
installation schedule. Our spare Leach controller is at CU for testing the
etalon.
TripleSpec has been operating with a borrowed Leach power supply so we can
monitor the array power dropouts.
Agile is still available at NA2.
APOLLO remains operational with some recent runs yielding significant lunar
returns. A gravimeter has been installed in the cone room beneath the
telescope, and it is producing useful positional information. The aircraft
transponder receiver array is also working quite well.
3) CIF projects
The direct-drive project is well on its way--nearly all components have been
ordered and drawings have been sent to the machine shops. The current plan is
to install these during the long engineering shutdown in August. The upgrade
project includes direct-drive motors with new axis controllers and encoders for
az and alt. We are also looking at replacing the current fiducial system with
a tape encoder, initially on the altitude axis.
**********************************
Imaging camera study:
Cynthia Froning presented the status of the study to upgrade our visible-light
imaging capabilities. Not many comments were received from a general call to
the community for input, so Cynthia has been in touch with SPIcam and other
users directly. There seems to be no obvious consensus on what upgraded
capabilities the users want, so the study is concentrating on two possible
designs with various options. Most users desire a larger field of view, which
can be as large as 7 to 9 arcminutes with the existing baffling at NA2 or the
Agile port. One of the options involves replacing the CCD and electronics with
a detector that has more, and smaller, pixels and lower read noise. The other
option is to provide multi-color simultaneous imaging capability. Cynthia will
produce a preliminary report this month that details the two design paths and
options, and a second report in June with cost and schedule projections.
Suzanne asked about the fast- guiding camera design used at Calar Alto, and
Cynthia said that she had looked at tip-tilt and "lucky" observing techniques,
but to implement fast-guiding would be expensive and would need to be justified
by a compelling specific science case; it is not clear that fast guiding is
desired for general imaging requirements, and the cost- benefit issue is a
problem. Jon Holtzman said he had looked at data from Fastcam which showed, in
simulations, that the potential benefit was modest, although the data were
taken in white light. Jon wondered what would be predicted if we were to use
comparable I-band images. Bill said that there is no filter holder in Fastcam,
but a filter could be mounted in front of the focal reducer if one wanted to
obtain these data.
Suzanne mentioned that if other users have additional opinions or different
ideas, they should speak up now before a decision is taken on a path forward.
**********************************
Echelle upgrade study:
Eric Burgh has approached the primary users of the echelle and has gotten
recommendations from several people. The key desired improvements are lower
read noise, better blue response, higher throughput and resolution, and smaller
pixels. The spectrograph has numerous vacuum-glass uncoated surfaces, and with
today's broadband coatings, an improvement in efficiency of ~28% is achievable,
according to Eric (but Jon Holtzman thought that earlier studies of this
predicted a nearly factor of two improvement; this needs more
investigation). The detector could be replaced with an e2v or Fairchild
device, yielding ~2 electrons read noise with 15-micron pixels. The smaller
pixels would also provide an improvement in resolution from the current value
of 33k to 54k, although it was pointed out that the slit-width constraints
might mitigate this gain. The mirror coatings are probably in good shape,
having lived in a soft vacuum all these years. The concept of replacing the
echelle grating with something more efficient is probably difficult due to
space constraints in the instrument. There was no strong user sentiment to
completely rebuild the echelle (e.g., split the red and blue into two channels,
add image slicers, etc.). There was some discussion of who within ARC has any
knowledge or memory of what the inside of the instrument looks like, and the
names of some people were passed on to Eric for future contacts. Bruce and Jon
asked what the RV performance level the echelle could be attainable with any or
all of these upgrades, and the answer is that this needs further study. Mark
suggested thinking about taking the echelle off the telescope and feeding it
with a fiber, which would enable use of the NA1 port by other instruments.
Suzanne said that the Users Committee members should canvas their users one
last time for input to this process. Eric should complete his proposal, and
encourage current and potential users to think harder about what upgrades to
the echelle would be compelling to make.
**********************************
User accounts on Newton:
Suzanne talked about the accounts that are set up on Newton for visiting APO
users to look at their data while on site. These
are institution-wide accounts (e.g. uwobserver, puobserver), but
these accounts sometimes are set up in a personal ad hoc way by
an on-site observer that makes it difficult for subsequent users and the APO
staff to use after that visitor leaves APO. She proposed that there should be
a "return-to-default" protocol employed when the visiting user is finished with
a user account during their visit. We could add more console-only logins, but
we don't want a large number of off-site accounts active on Newton. Jon
recommended that users who visit the site often could bring their own temporary
setup scripts. We are
looking into putting a policy in place to make these user accounts return to a
default state.
**********************************
Telescope scheduling process:
Suzanne and Russet sent a write-up of the 3.5-m scheduling process to the User
Committee members, and asked the committee members to read the screed and
advise if it is ready to send to the general community and be posted. She will
send it to the institutional schedulers in any case, and eventually it will be
posted on the scheduling web page.
**********************************
Miscellaneous [added in proof]:
Suzanne forgot to mention that Jon Holtzman submitted a white paper to the
Decadal Survey, titled "Apache Point Observatory: Facilities, Operations, and
Partnerships", which can be downloaded at
http://www7.nationalacademies.org/bpa/Astro2010_Position_Papers_byTitle.html
Suzanne thanks Jon and the other authors for putting this together and getting
it posted.
**********************************
ACTION ITEMS:
[open from previous months]:
==>ACTION: Suzanne and Russet to write a draft description of our scheduling
process for eventual user dissemination.
STATUS: Completed, closed this meeting.
[new actions from this meeting]:
none
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 11 May, at
8:30 AM Pacific Time. The agenda and other materials will be sent to the
committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1130 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO